### v2

# ASME Letters in Translational Robotics

[…]
pdfstandard=A-3u,pdfversion=1.7, lang=en-US,
[…]
ASME Letters in Translational Robotics
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
email: agung@ecse.rpi.edu
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
[…]
Rensselaer Polytechnic Institute,
Troy, NY 12180 USA
email: wenj@rpi.edu
Multi-Robot Scan-n-Print for Wire Arc Additive Manufacturing
[…]
Zusammenfassung
Robotic Wire Arc Additive Manufacturing (WAAM) is a metal additive manufacturing technology offering flexible 3D printing while ensuring high-quality near-net-shape final parts. However, WAAM also suffers from geometric imprecision, especially for low-melting-point metal such as aluminum alloys. In this paper, we present a multi-robot framework for WAAM process monitoring and control. We consider a three-robot setup: a 6-dof welding robot, a 2-dof trunnion platform, and a 6-dof sensing robot with a wrist-mounted laser line scanner measuring the printed part height profile. The welding parameters, including the wire feed rate, are held constant based on the materials used, so the control input is the robot path speed. The measured output is the part height profile. The planning phase decomposes the target shape into slices of uniform height. During runtime, the sensing robot scans each printed layer, and the robot path speed for the next layer is adjusted based on the deviation from the desired profile. The adjustment is based on an identified model correlating the path speed to changes in height. The control architecture coordinates the synchronous motion and data acquisition between all robots and sensors. Using a three-robot WAAM testbed, we demonstrate significant improvements of the closed loop scan-n-print approach over the current open loop result on both a flat wall and a more complex turbine blade shape.
keywords:
WAAM, Welding Robot, Multiple Robot, Scan-n-Plan, Welding Process Control
1 INTRODUCTION
Wire Arc Additive Manufacturing (WAAM) is a metal additive manufacturing technology that builds layers of a target shape by melting metal wire using an electric arc, often in conjunction with a motion system [1]. By combining the welding system with an industry robot, a robotic WAAM system can produce large format near-net-shape parts at high deposition rate. It is ideally suited for rapid metal prototyping and repairs of forged parts in aerospace and automotive industries [2].
Abbildung 1: Scan-n-Print WAAM testbed
Robotic WAAM processes present several challenges, with the foremost being geometry imprecision, which compromises near-net-shape capability, especially when working with low-melting-point metals such as aluminum alloys [3]. When operating commercial WAAM systems in the open-loop mode, defects can arise intermittently. Addressing these variations is essential to prevent error accumulation, which can compromise manufacturing quality. Some methods model the process using Gaussian Process Regression Models [4, 5] and apply reinforcement learning for optimization. Others incorporate additional control actions, such as arc control combined with milling [6], or use milling strategies to mitigate error accumulation [7]. However, these approaches are less efficient and increase material usage because of the post-processing as compared to direct process improvement. Many of these methods rely on laser scanners to capture weld geometry profiles, while others utilize alternative data sources, such as acoustic signals [8, 9], for defect detection, enabling real-time correction and improved efficiency.
A robotic WAAM system typically involves a 6-dof welding robot combined with a 2-dof trunnion to enable alignment of metal deposition with gravity [10]. An addition 6-dof robot mounted with sensors provides comprehensive inspection coverage. Motion of these robotic devices need to be coordinated to ensure task execution and avoidance of constraints such as joint limits and collision. Planning techniques to avoid robot singularities [11] and selection of welding paths [12] are commonly employed. The standard approach involves preplanned paths with consideration of robot kinematics [13, 14]. Although many methods incorporate redundancy for added flexibility, effective strategies are needed to resolve it [15, 16]. To address these challenges, we consider a three-robot setup, including a 6-dof welding robot, a positioner with 2-dof trunnion platform and a 6-dof sensing robot with a wrist-mounted laser line scanning measuring the printed height profile. We use the motion planning approach for multiple robot arm from our previous work [17, 18] for robot arms path planning and [3] gravity alignment.
This paper presents a multi-robot framework for WAAM process monitoring and control. In the planning phase, the welding and inspection motion between the robots and the positioner are generated to meet the welding and inspection requirements. During the welding phase, we consider the robot torch path speed as the control input while holding the welding parameters, such as the wire feed rate at constant based on the materials used. The inspection robot measurements the layer height using a wrist-mounted laser scanner. The scans from the laser scanner are registered using the calibrated robot kinematics, followed by removal of outlier points to generate the height profile of the deposit. The torch speed for the next layer is then adjusted by the deviation from the desired profile. The adjustment is based on an identified model correlating the torch speed to the change in deposition height. The control architecture, described in [19], coordinates the motion and data acquisition between all robots and sensors. The closed loop WAAM framework has been applied in our testbed for printing a vertical wall, a complex turbine-blade-like geometry and a cylindrical geometry, using an aluminum alloy. The geometric fidelity of these parts outperform the open-loop baselines in both layer-by-layer and continuous-scan modes. We also apply our approach to a steel alloy, known for its higher melting point and greater thermal stability compared to aluminum alloys. The proposed techniques works well and shows improvements over the open loop printing. We demonstrate the repeatability of the manufacturing process, so that once the corrected motion profile is learned, the scanning step may be removed, thereby enhancing manufacturing efficiency. This paper is organized to present the scan-n-print framework in Section 2, the deposition model in Section 3, and experimental results and evaluation in Section 4. The source code, datasets and video related to this study are publicly available. [^1] https://github.com/rpiRobotics/Convergent_Manufacturing_WAAM/tree/main/scan_n_print  [^2] https://youtu.be/Ei2e2mBJZ5s
2 Scan-n-Print Framework
2.1 Overview
The proposed closed-loop WAAM framework is illustrated in Figure 2. After the deposition of each layer and scanning, the part geometry is reconstructed, with each scan registered to the positioner frame using the forward kinematics of the scanning robot and the positioner. Sampling the surface along the welding path gives the height profile of the top surface. For each layer, the slicing algorithm generates the desired layer height $\Delta h_{d}$. The measured height profile of the current layer is compared with the desired profile and the difference, together with the torch speed to height model, generates the robot motion for the next layer.
Abbildung 2: Scan-n-Print framework
2.2 Motion Generation and Execution
Motion planning is required for both the WAAM system (6-dof welding robot and 2-dof trunnion) and the sensing robot. As described in [3], motion of the combined 8-dof system is determined by the torch position based on the slicing of the target geometry, the commanded path velocity, and the alignment requirement of the gravity direction with the previous layer. There is one redundant degree of freedom for the rolling motion along the torch. This redundancy resolved using resolved motion control subject to speed and joint limit constraints. The scanning motion of the sensing robot is designed to ensure complete coverage of the top layer. As illustrated in Figure 3, the scan path is along the known welding trajectory while maintaining the appropriate laser-scan standoff distance. To fully cover critical areas at the start and end of the weld, side scanning motion is also used. As for the welding motion, the scanning path is designed in the workpiece frame and then transformed to the positioner frame once the workpiece is placed on the positioner. The robot motion is designed to ensure that the scans plane is perpendicular to the deposited layer. Since the gravity alignment is not needed, there is two redundant degrees of freedom (6-dof scanning robot and the 2-dof positioner). The same resolved motion control is used for redundancy resolution.
Abbildung 3: Scanning path plan
The overall system architecture to integrate multi-robot control and multi-sensor data acquisition is described in [3]. In the step-wise mode, the Cartesian motion of both the welding and scanning robots are segmented into uniform linear segments and encoded in the Motoman robot programming language, INFORM. The robot program together with the corresponding welding commands are sent to the Motoman DX200 robot controller. The DX200 controller coordinates the motion of the welding and scanning robots and trunnion table, and the Fronius weld controller. In the continuous mode, we bypass the Motoman controller and directly command the robot motion and weld controller. This is accomplished with two custom drivers using Robot Raconteur (RR) as the middleware [20]. The robot driver is developed using the Motoman MotoPlus package to stream robot motion setpoints and read robot joint angles at 8 ms to all three robots. The weld controller driver interfaces to the Fronius weld controller to change weld parameters and turn the welding arc on and off at 10 Hz.
2.3 Deposition Height Profile
During run time, each layer is scanned immediately after printing. The laser scans form a 3D point cloud in the laser scanner frame. To transform the points to the robot frame, we used the robot joint angles and the inverse kinematics which requires robot kinematic calibration [21]. To mitigate noise, we apply statistical and cluster-based outlier removal, and a average smoother along the welding direction. The cleaned point cloud enables extraction of the weld top profile by sampling along welding travel direction. With a well-defined point cloud representation of the weld, we extract the layer deposition profile by slicing the point cloud along the printing direction. The height of each location is determined by the height of the points within the point cloud. Figure 4 displays the top layer height profile of the welded piece in the $x$-axis direction.
Abbildung 4: Deposition height of the weld piece
2.4 Motion Optimization
The motion optimization module aims to achieve uniform deposition height and correct welding defects. It is an essential step to prevent error accumulation and ensure consistent surface quality. We adjust only the torch path speed while maintaining constant wire feed rate to regulate the amount of deposited materials at each time. Denote relationship between the torch speed and the weld deposition height at a given feed rate $\Delta h=f(v)$, where $v$ is the torch speed and $\Delta h$ the deposition height. Empirical characterization of this model is described in Section 3. In general, $f$ is a monotonically decreasing function as higher speed means less time for deposition and therefore smaller $\Delta h$.
Denote the average height of the current layer, $h^{(i)}$, calculated from the deposition profile. The target height for the next layer, $h^{(i+1)}_{d}$, is $h^{(i+1)}_{d}=h^{(i)}+\Delta h_{d}^{(i)}$, where $\Delta h_{d}^{(i)}$ is the desired deposition height. The robot motion for each layer is divided into linear torch motion segments. Let the average measured height for the $k$th segment be $\bar{h}^{(i,k)}$. The desired deposition height for this segment is then calculated as $\Delta h^{(i,k)}_{d}=h^{(i+1)}_{d}-\bar{h}^{(i,k)}$. Since the deposition model is monotonic, it is invertible. The inverse model then provides the required path speed:
$v^{(i,k)}=f^{-1}\left(\Delta h^{(i,k)}_{d}\right).$ (1)
Figure 5 shows the measured heights and the corresponding commanding speeds for the motion segments along a typical layer. We apply a moving average smoothing filter to the height profile. This is a low pass filter that remove high frequency components in the profile that could cause unwanted high acceleration. The maximum and minimum speeds are also capped to prevent exaggerated movements. If a desired speed is not achieved because of the limited acceleration and speed, the remaining height errors are corrected in the subsequent layers. The optimized motion speed is commanded in the robot motion program if operated in the step-wise mode. In the continuous mode, we directly update the torch speed by adjusting the robot joint trajectory streamed to the robot controller in real time.
Abbildung 5: Speed update of 40 motion segments. Dots: difference from the current height to the target height $\Delta h_{d}$. Lines: the updated torch speed.
2.5 Step-wise vs Continuous Scan-n-Print
The scan-n-print framework operates in step-wise or continuous modes. In the step-wise mode, robots alternate between welding and scanning, with the deposition profile processing and motion optimization occurring after the scanning motion. In the continuous mode, the scanning robot provides real-time feedback by looking ahead of the layer height, allowing on-the-fly processing and optimization. The continuous mode offers more immediate feedback and correction, and does not require turning the welding arc on and off (which compromises deposition uniformity) between layers. However, the implementation of the continuous mode is more complex due to the coordinated motion between all three robots The motion optimization method is the same in both modes, but the update frequency is different. The step-wise mode optimizes the motion of the entire layer, while the continuous mode optimizes the motion of the current welding point during run-time using the look-ahead scanner. In our setup, the continuous mode closes the feedback loop at 100 Hz, limited by the scanning rate. Both modes produce high quality final parts. We observe that the continuous mode has slightly better performance in terms of the height standard deviation. This is because the step-wise mode requires turning the welding arc on and off at the edges.
3 Welding Model
To derive the model linking welding motion speed to deposition height using the collected layer data, we reference the deposition equation in [22]:
$S_{w}v_{MR}=S_{B}v$ (2)
where $S_{w}$ is the wire cross-section, $v_{MR}$ is the wire melting rate, $S_{B}$ is the weld bead cross-section, and $v$ is the welding torch speed. Note that $v_{MR}$ is proportional to the wire feed rate. Assuming that the cross-section of the weld bead width is proportional to the height with proportionality constant $c$, we have approximately
$\frac{S_{w}v_{MR}}{c\Delta h^{2}}=v.$ (3)
It follows that
$\ln(\Delta h)=-\frac{1}{2}\ln(v)+\frac{1}{2}\ln\left(\frac{S_{w}v_{MR}}{c}\right)$ (4)
To account for inaccuracy in the model, including the complex weld shape, we consider a more general expression:
$\ln(\Delta h)=a\ln(v)+b$ (5)
where $a$ and $b$ are constants that depend on the wire feed rate, to be determined based on collected data.
To identify the model, we collect data as follows:
  • Weld two base layers for a consistent foundation for each run.
  • Reduce the torch speed by 2mm/sec every two layers, from 20mm/sec to 2mm/sec, while maintaining a constant feed rate.
  • Measure the average height of each welded layer.
Figure 6 shows the layer heights of a weld sample collected under fixed wire feed rate at 100 ipm (inch per minute). As expected, the graph shows that an increase in torch speed results in a decrease in deposition height. Figure 7 presents the identified model parameters, with least square fit of a line to the collected data. By repeating this procedure with varying feed rates, model functions for different wire feed rates can be developed.
Abbildung 6: 100 ipm layers height. The speed of the torch varies from 2 mm/sec at the bottom layers to 20 mm/sec at the top layers to collect deposition data at each traveling speed.
Abbildung 7: Welding motion speed to deposition model for 100 ipm wire feed rate. Table. 1 shows the model parameters of all tested wire feed rate with the data fitting RMSE.
The identified coefficients are provided in Table 1. Note that the coefficients of the first-order terms are consistently very close to $-0.5$ which aligns with the equation derivation. We also showed the RMSE of the data after fitting the model. During the scan-n-print operation, the wire feed rate is fixed at a chosen value and the torch speed is adjusted based on the scanning result.
feed rates 100 110 130 150 170 240
a (slope) -0.62 -0.43 -0.43 -0.44 -0.45 -0.46
b 1.85 1.23 1.630 1.37 1.40 1.15
RMSE 0.27 0.12 0.13 0.18 0.21 0.10
Tabelle 1: Identified model parameters and regression error (in mm) under various constant wire feed rates (in inch-per-minute)
4 Experiments and Evaluation
4.1 Hardware Testbed
The hardware testbed is shown in Figure 1. It comprises a MA2010 welding robot with a D500B positioner, and a MA1440 inspection scanning robot, all from Yaskawa Motoman. Fronius 500i provides the welding power source for the cold metal transfer process. The Fronius weld controller is integrated with the robot controller DX200. The weld material is ER 4043, an aluminum alloy with lower melting point compared to materials like steel alloy or stainless steel. Its high thermal conductivity often leads to varying weld conditions and a higher occurrence of defects. In Section 4.4, we apply scan-n-print to ER 70S-6, a steel alloy with higher melting point and lower thermal conductivity than the aluminium alloy. For laser scanning, we use an MTI Pro-Track G Series laser scanner installed on the inspection robot. The scanner must maintain a standoff distance between 65mm to 125mm with a field of view of approximately 40mm to 60mm. The scans are collected at a frame rate of about 100 Hz.
Abbildung 8: Robotics WAAM and scanning execution
4.2 Performance Comparison: Wall Structure
For the baseline, we select the best combination of wire feed rate and torch speed: 100 ipm and 5mm/sec. With the identified model in Section 3, the input height is approximately $\Delta h_{d}\approx 2.34$mm when the torch speed is set to 5mm/sec. For comparison between the scan-n-print and open loop, we print a wall structure with width $65$mm and thickness $4$mm, and the top layer exceeding 50mm in height.
(a) Open-loop Baseline
(b) Scan-n-Print framework
Abbildung 9: Open-loop baseline vs. Scan-n-Print for a Wall Structure, ER 4043 Aluminum Alloy.
Figure 9 shows the printed pieces produced for the baseline and scan-n-print. Even through visual observation alone, it is evident that the quality of the printed piece has significantly better layer uniformity with the closed-loop approach. Note that defects tend to occur predominantly at the edges of the weld piece due to the arc on/off effects. Figure 10(a) shows the standard deviation of the layer height for both the open-loop baseline and closed-loop correction tests. In the open loop case, the height variation in each layer accumulates with the number of layers. To evaluate the print performance without the edge effects, we exclude the edge data. We define the edges as within 7.5757.57.5mm from the welding arc on and off points. We can identify the regions on the point cloud using the collected robot joints data and the associated scans. Figure 10(b) shows that the closed-loop approach still outperforms the open-loop approach. Table 2 shows the statistics and improvements of the height uniformity over all layers. For the wall geometry, the standard deviation of height improved 66%. Even without the edges where defects dominate, the performance improved by 60%.
Mean Height STD With edge Without edge
Baseline (mm) 1.38 0.33
Correction (mm) 0.47 0.13
Improvement (%) 66 60
Tabelle 2: Mean height standard deviations across all layers, with and without the edges, of the WAAM wall, for open-loop baseline and scan-n-print correction cases, ER 4043 Aluminum Alloy.
(a) Full piece
(b) Edge excluded
Abbildung 10: Height STD of each layer of the WAAM wall. Open-loop baseline vs closed-loop correction, ER 4043 Aluminum Alloy.
We further illustrate the accuracy comparison of the printed WAAM wall structure versus the desired CAD model as shown in Figure 11(a). The WAAM part is scanned using the Artec Spider scanner to generate a point cloud representation. It is then aligned with the CAD model using Iterative Closest Point (ICP) registration [23]. The accuracy is characterized by the Euclidean distance error between the printed WAAM part and the desired structure. Figure 12 shows the error heat map on the wall CAD model. The heat map reinforces the observation that the proposed framework is far more accurate than the baseline. The error is mostly due to the defect on the edges. Table 3 shows the error statistics. The maximum error improves by 70% and the average error improves by 13%.
(a) Wall 3D profile
(b) Blade 3D profile
Abbildung 11: The 3D profile of the printed wall and blade (point cloud data from the Artec scanner) vs. the desired structure CAD model (blue).
(a) Baseline
(b) Correction
Abbildung 12: Error heat-map between the desired wall structure CAD model and the WAAM wall.
Wall Average Error Max error
Baseline (mm) 0.39 5.38
Correction (mm) 0.34 1.52
Improvement (%) 13 72
Tabelle 3: Average and maximum distance between the WAAM wall and the CAD model
4.3 Performance Comparison: Blade-like Structure
For a more complex geometry, we apply the proposed framework to a generic fan blade geometry. The experiment was conducted following the same setup as the wall structure, with the wire feed rate at 100 ipm and desired deposition height of $\Delta h_{d}=2.34$mm.
(a) Open-loop WAAM blade
(b) Scan-n-Print WAAM blade
Abbildung 13: A generic fan blade geometry using the proposed scan-n-print framework. The red circle shows the defect when using the open-loop setting.
Figure 13 shows the blade both in open loop and scan-n-print modes. A visual comparison reveals a clear defect at the edge (indicated by the red circle) in the open loop case, where the welding torch fails to deposit sufficient material at the edge of the piece. The closed-loop print shows a far smoother edge. Figure 14(a) shows the standard deviation and root mean square error (RMSE) of the deposition heights at each layer. The larger error at the top layers is due to the short segments where the smaller number of measurements tend to accentuate these metrics. Table 4 shows the statistics and the improvements. The height STD improved by 53 % and tracking rmse improved by 45 %.
(a) Height STD
(b) Height RMSE
Abbildung 14: Height STD and RMSE of each layer of the WAAM blade
Blade Mean Height STD Mean Tracking RMSE
Baseline (mm) 0.85 0.90
Correction (mm) 0.24 0.49
Improvement (%) 53 45
Tabelle 4: Mean height STD and tracking RMSE across the printed layers of the WAAM blade.
The comparison of the printed blade and the blade CAD model is shown in Figure 11(b). Table 5 summarizes the error statistics. The maximum error improves by 5% and the average error improves by 13%.
Blade Average Error Max error
Baseline (mm) 0.75 2.23
Correction (mm) 0.65 2.11
Improvement (%) 13 5
Tabelle 5: Average and maximum distance between the WAAM blade and the CAD model
4.4 Performance Comparison: Steel Alloy
We follow the same procedure in Section 4.2 but using the steel alloy ER 70S-6. We select the best combination of wire feed rate and torch speed for the steel alloy: 200 ipm and 7 mm/sec. We apply the same model identification procedure in Section 3. The input height is approximately $\Delta h_{d}\approx 1.35$mm when the torch speed is set to 7mm/sec. The wall structure has the same dimension with width 65 mm and thickness 4 mm.
(a) Open-loop Baseline
(b) Scan-n-Print framework
Abbildung 15: Open-loop baseline vs. Scan-n-Print for a Wall Structure, ER 70S-6 Steel Alloy
Figure 15 shows the printed steel pieces for the baseline and scan-n-print. By visual observation, the open-loop welding with steel alloy does not create as much edge defects as with the aluminum alloy. This is because the steel alloy has higher melting point and lower thermal conductivity and is more stable than aluminum alloy. However, the top layer is still slightly curved compare to the close-loop approach. Figure 16(a) shows the standard deviation of the layer height for both the open-loop baseline and closed-loop correction tests. The open-loop case shows the error accumulation, but not as much as in the aluminum alloy case. The proposed scan-n-print still shows improvement. We exclude the edge data and evaluate the performance without edge effects. Figure 16(b) shows that the uniformity between the open-loop and the closed-loop approach is comparable. However, we see a trend of error accumulation in the open-loop case at the end of the weld and the height STD exceeds the closed-loop case significantly. Table 6 shows the statistics and improvements of the steel alloy case. As expected, the steel alloy is stable for WAAM so the mean height does not vary as much as the aluminum alloy in the open-loop baseline case (1.47 vs 0.74) but the proposed scan-n-print can further improve the weld by 26%. Without considering the edges, the performance improves by 10%.
Mean Height STD With edge Without edge
Baseline (mm) 0.74 0.31
Correction (mm) 0.55 0.28
Improvement 26% 10%
Tabelle 6: Mean height standard deviations across all layers, with and without the edges, of the WAAM wall, for open-loop baseline and scan-n-print correction cases, ER 70S-6 Steel Alloy.
(a) Full piece
(b) Edge excluded
Abbildung 16: Height STD of each layer. Open-loop baseline vs closed-loop correction, ER 70S-6 steel alloy.
4.5 Continuous Scan-n-Print
We also demonstrated the scan-and-print framework in a continuous mode setup, with the same control algorithm, targeting a cylindrical geometry with a radius of $r=35$ mm. In this setup, the scanning robot is positioned half the perimeter ahead of the welding torch. The leading scanner provides look-ahead feedback by capturing the coming geometry surface and storing it in the memory. The torch speed is then adjusted while reaching the corresponding position. The feed rate was set at 160 ipm with a desired deposition height of $\Delta h_{d}=1.80$ mm. Figure 17 shows the results of the continuous scanning and printing process. The significant reduction in standard deviation demonstrates the effectiveness of the proposed framework in smoothing the height variation in each layer.
(a) Cylinder
(b) Height STD
(c) Height RMSE
Abbildung 17: Continuous Scan-n-Print of a cylinder.
4.6 Repeatability Test
For printing of multiple parts, we may use scan-n-print to determine the torch speed in each layer and then apply the recorded torch speed profiles in the more efficient open loop printing of subsequent parts. Figure 18 shows the comparison between the part produced by the scan-n-print and parts produced by replaying the recorded velocity commands. Qualitatively, the replayed printing match closely the original closed loop print, with the notable absence of the edge defects observed in the baseline part. The top layer also appears to be relatively uniform. The standard deviation of layer heights, shown in Figure 19(a), corroborates these observations. When the edge regions are excluded from the analysis, shown in Figure 19(b), the replayed prints perform similarly as the open loop, indicating that the correction is mostly for the edge of the wall.
(a) Recorded
(b) Repeat 1
(c) Repeat 2
Abbildung 18: Repeatability test.
(a) Full Piece
(b) Edge Excluded
Abbildung 19: Repeatability test: height STD of each layer. Open-loop baseline vs closed-loop approach vs repeats
5 Conclusion
In this paper, we introduced a scan-n-print framework to improve the performance of the WAAM processes. The height variation is measured using a laser scanner and the torch speed is adjusted to compensate for the height variation. Layer height regulation is accomplished by inverting an identified model relating the torch speed to the deposition height. The effectiveness of the proposed approach is demonstrated on several printed geometries, vertical wall, a mock fan blade, and a cylinder. The demonstration is conducted on our WAAM testbed with a welder robot and positioner, and a separate monitoring robot with a mounted laser scanner. We are currently working on regulating both the geometry and temperature by controlling both wire feed rate and torch speed and using an IR camera for temperature feedback in addition to the laser scanner.
ACKNOWLEDGMENT
The authors would also like to thank Chris Anderson and Roger Chistian at Yaskawa Motoman, Jeff Schoonover at GE Aerospace, and Matt Robinson at Southwest Research Institute for their helpful discussion of the project and Terry Zhang for his help with the WAAM hardware.
Funding Data
Research was sponsored by the ARM (Advanced Robotics for Manufacturing) Institute through a grant from the Office of the Secretary of Defense and was accomplished under Agreement Number W911NF-17-3-0004. The views and conclusions con- tained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Office of the Secretary of Defense or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Literatur
  • [1] Williams, S. W., Martina, F., Addison, A. C., Ding, J., Pardal, G., and Colegrove, P., 2016, “Wire + Arc Additive Manufacturing,” Materials Science and Technology, 32(7), pp. 641–647.
  • [2] Busachi, A., Erkoyuncu, J., Colegrove, P., Martina, F., and Ding, J., 2015, “Designing a WAAM Based Manufacturing System for Defence Applications,” Procedia CIRP, 37, pp. 48–53, CIRPe 2015 - Understanding the life cycle implications of manufacturing.
  • [3] He, H., lung Lu, C., Ren, J., Dhar, J., Saunders, G., Wason, J., Samuel, J., Julius, A., and Wen, J. T., 2024, “Open-Source Software Architecture for Multi-Robot Wire Arc Additive Manufacturing (WAAM),” arxiv:2408.04677, https://arxiv.org/abs/2408.04677
  • [4] Xiong, Y., Park, S.-i., Padmanathan, S., Dharmawan, A., Foong, S., Rosen, D., and Soh, G., 2019, “Process planning for adaptive contour parallel toolpath in additive manufacturing with variable bead width,” International Journal of Advanced Manufacturing Technology, 105, pp. 4159–4170.
  • [5] Dharmawan, A. G., Xiong, Y., Foong, S., and Song Soh, G., 2020, “A Model-Based Reinforcement Learning and Correction Framework for Process Control of Robotic Wire Arc Additive Manufacturing,” 2020 IEEE International Conference on Robotics and Automation (ICRA), Paris, France, May 31 –- June 4, pp. 4030–4036, 10.1109/ICRA40945.2020.9197222.
  • [6] Ma, G., Zhao, G., Li, Z., Yang, M., and Xiao, W., 2019, “Optimization strategies for robotic additive and subtractive manufacturing of large and high thin-walled aluminum structures,” The International Journal of Advanced Manufacturing Technology, 101, pp. 1275–1292.
  • [7] Lim, W. S., Dharmawan, A. G., and Soh, G. S., 2022, “Development and performance evaluation of a hybrid-wire arc additive manufacturing system based on robot manipulators,” Materials Today: Proceedings, 70, pp. 587–592, The International Conference on Additive Manufacturing for a Better World (AMBW 2022).
  • [8] 2021, A Study on the Acoustic Signal Based Frameworks for the Real-Time Identification of Geometrically Defective Wire Arc Bead, Vol. Volume 3A: 47th Design Automation Conference (DAC) of International Design Engineering Technical Conferences and Computers and Information in Engineering Conference, Online, Virtual, August 17 –- 20.
  • [9] Surovi, N. A. and Soh, G. S., 2023, “Acoustic feature based geometric defect identification in wire arc additive manufacturing,” Virtual and Physical Prototyping, 18(1), p. e2210553.
  • [10] Kim, D.-W., Choi, J.-S., and Nnaji, B. O., 1998, “Robot arc welding operations planning with a rotating/tilting positioner,” International Journal of Production Research, 36(4), pp. 957–979.
  • [11] Ahmad, S. and Luo, S., 1989, “Coordinated motion control of multiple robotic devices for welding and redundancy coordination through constrained optimization in Cartesian space,” IEEE Transactions on Robotics and Automation, 5(4), pp. 409–417.
  • [12] Bhatt, P. M., Kulkarni, A., Malhan, R. K., and Gupta, S. K., 2021, “Optimizing Part Placement for Improving Accuracy of Robot-Based Additive Manufacturing,” 2021 IEEE International Conference on Robotics and Automation (ICRA), Xi’an, China, May 30 -– June 5, pp. 859–865, 10.1109/ICRA48506.2021.9561494.
  • [13] Schmitz, M., Wiartalla, J., Gelfgren, M., Mann, S., Corves, B., and Hüsing, M., 2021, “A Robot-Centered Path-Planning Algorithm for Multidirectional Additive Manufacturing for WAAM Processes and Pure Object Manipulation,” Applied Sciences, 11(13), p. 5759.
  • [14] Michel, F., Lockett, H., Ding, J., Martina, F., Marinelli, G., and Williams, S., 2019, “A modular path planning solution for Wire + Arc Additive Manufacturing,” Robotics and Computer-Integrated Manufacturing, 60, pp. 1–11.
  • [15] Lizarralde, N., Coutinho, F., and Lizarralde, F., 2022, “Online Coordinated Motion Control of a Redundant Robotic Wire Arc Additive Manufacturing System,” IEEE Robotics and Automation Letters, 7(4), pp. 9675–9682.
  • [16] Hu, Y., Huang, B., and Yang, G.-Z., 2015, “Task-priority redundancy resolution for co-operative control under task conflicts and joint constraints,” 2015 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS), Hamburg, Germany, September 28 -– October 2, pp. 2398–2405, 10.1109/IROS.2015.7353702.
  • [17] He, H., Lu, C.-L., Wen, Y., Saunders, G., Yang, P., Schoonover, J., Wason, J., Julius, A., and Wen, J. T., 2023, “High-Speed High-Accuracy Spatial Curve Tracking Using Motion Primitives in Industrial Robots,” 2023 IEEE International Conference on Robotics and Automation (ICRA), London, United Kingdom, May 29 –- June 2, pp. 12289–12295, 10.1109/ICRA48891.2023.10161022.
  • [18] He, H., Lu, C.-L., Saunders, G., Wason, J., Yang, P., Schoonover, J., Ajdelsztajn, L., Paternain, S., Julius, A., and Wen, J. T., 2024, “Fast and Accurate Relative Motion Tracking for Dual Industrial Robots,” IEEE Robotics and Automation Letters, 9(11), pp. 10153–10160.
  • [19] He, H., Aksoy, B., Saunders, G., Wason, J., and Wen, J. T., 2023, “Plug-and-Play Software Architecture for Coordinating Multiple Industrial Robots and Sensors from Multiple Vendors,” 2023 IEEE 19th International Conference on Automation Science and Engineering (CASE), Auckland, New Zealand, August 26 – 30, pp. 1–8, 10.1109/CASE56687.2023.10260612.
  • [20] Wason, J. D. and Wen, J. T., 2023, “Robot Raconteur®: Updates on an Open Source Interoperable Middleware for Robotics,” 2023 IEEE 19th International Conference on Automation Science and Engineering (CASE), Auckland, New Zealand, August 26 – 30, pp. 1–8, 10.1109/CASE56687.2023.10260569.
  • [21] Xu, J., Hoo, J. L., Dritsas, S., and Fernandez, J. G., 2022, “Hand-eye calibration for 2D laser profile scanners using straight edges of common objects,” Robotics and Computer-Integrated Manufacturing, 73, p. 102221.
  • [22] Handa, H., Okumura, S., and Nio, S., 1997, “The robotic easy teaching system in computer aided welding,” NIST Special Publication(USA),, 923, pp. 562–575.
  • [23] Besl, P. and McKay, N. D., 1992, “A method for registration of 3-D shapes,” IEEE Transactions on Pattern Analysis and Machine Intelligence, 14(2), pp. 239–256.